OpenStudy (anonymous):

Will give medal, please help. Simplify:square root of 5 open parentheses 8 plus 3 square root of 6 close parentheses

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Can you use the equation tool to make it easier to read?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

\[7\sqrt{3}-4\sqrt{6}+\sqrt{48}-\sqrt{54}\]

OpenStudy (anonymous):

\[11\sqrt{3}-\sqrt{6}\]

OpenStudy (anonymous):

Thanks :) Could you do somemore please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Thanks I have 3 more.... Is that fine?

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

\[\sqrt{10} * \sqrt{8}\]

OpenStudy (anonymous):

\[4\sqrt{5}\]

OpenStudy (anonymous):

Wow Thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[5\sqrt{7}+12\sqrt{6}-10\sqrt{7}-5\sqrt{6}\]

OpenStudy (anonymous):

\[-5\sqrt{7}+7\sqrt{6}\]

OpenStudy (anonymous):

ok thanks here somemore

OpenStudy (anonymous):

\[4\sqrt{2}+7\sqrt{2}-3\sqrt{2}\]

OpenStudy (anonymous):

8 what?

OpenStudy (anonymous):

\[8\sqrt{2}\]

OpenStudy (anonymous):

ok 1 more

OpenStudy (anonymous):

Is the simplified form of \[2\sqrt{3} - 2\sqrt{3}\] rational. Yes or No?

OpenStudy (anonymous):

yes, because it equals to 0

OpenStudy (anonymous):

Ok.... Thank you so much for your help. :)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

are you available for more. Im like so sorry more came up.?

OpenStudy (anonymous):

@sujuexob2st